To: J. Cook, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IR Demo Project Weekly Report, May 4-8, 1998
Date: May 8, 1998
Commissioning activities continued all week on the FEL except
for a single shift of scheduled maintenance. Highlights for the
week include: (1) successful operation of the M55 diagnostic
for measuring phase and bunch length and (2) check-out of the
optical control room diagnostics.
Progress was made in defining a JLab collaboration with the National
Renewable Energy Laboratory (NREL) for use of the FEL for characterization
and processing of photovoltaic materials. F. Dylla and L. Kazmerski
(NREL) met last weekend to draft a proposal that will be sent
to both the DOE-BES and EE program offices.
On Monday, May 4, the City of Newport News and Jefferson Lab hosted
the official dedication of the Applied Research Center (ARC) building.
Over 400 people attended the event which was highlighted by official
remarks by Jim Gilmore, Governor of Virginia, and Barry Duval,
State Secretary of Commerce and Trade. JLab and members of ARC
university consortium presented 20 posters on the building that
highlighted R&D activities that will be moved into the building.
Many of those activities have close linkage to proposed FEL user
The last DY 180° bend magnet was assembled with its vacuum
chamber and is ready for installation later this month.
The 2 skew quads around the cryomodule were installed and tested
this week. They are now operational.
A strong back for the wiggler was completed this week. It is
ready to install the wiggler later this month.
Phase stability tests were conducted this week to determine the
phase relationship between the drive laser and the buncher. Although
several problems were uncovered, no direct link was established
to the 10° phase drifts we have been experiencing.
All of the BPMs in the 4F and 5F regions were tested this week
for broken sensors. Only one BPM had any broken sensors. The
faulty 2F02 BPM is being replaced this weekend.
A new RF 4 seater combo screen was installed this week and is
now available for the injector. A revised fault screen for the
injector was built and will be available early next week.
This week's activities centered on continued commissioning of
diagnostics, including development of procedures to use them,
additional difference-orbit measurements, and developing a reproducible
injector-setup procedure that involves diagnostics downstream
of the injector. None of these activities are complete yet, but
progress was made in all of them. In addition, one shift was
used to devise a BLM setup that would permit operating the Happek
interferometric bunch-length monitors as well as to recheck the
BLM settings for cw running. Regarding the latter, the voltage
of the BLM at the straight-ahead dump was adjusted, but otherwise
the existing setup looked appropriate.
One complication is a persistent drift in the rf phase of the
drive laser that began to be observable about two weeks ago.
The drift causes the beam properties to change concomitantly,
and it must be eliminated prior to trying to lase. We believe
that it was introduced as a consequence of some hardware changes
in the drive-laser rf system, and we have been trying to track
down the cause using well-defined test plans. At this writing
(11:15, 8 May 98) we still had not done so definitively, but we
have suspected causes that we will check over the next few days.
The status of the electron-beam diagnostics is as follows:
- MaxVideo works for beam size measurements. However, on several of the viewers noise and dead pixels prevent its use because the signal-to-noise ratio is too low.
- Current monitor cavity works and is in constant use.
- BPMs work; 2 cans (1 injection line, 1 linac) are bad but can reproduce orbits. One of the bad cans is the key BPM in front of the wiggler, and it needs to be replaced to enable setup for lasing. Replacement is scheduled to occur tomorrow (Saturday). We don't yet know whether the BPMs have the resolution or stability needed for FEL setup.
- Multislits work, including analysis software; we are deciding on the best way to use them. Neither of the multislits presently gives an emittance measurement in the standard beam setup. It may be possible to get the needed measurements by altering quad settings and then backing out the numbers, and we are presently working on the methodology for doing so.
- Quad/viewer works; we need the final setup for good numbers. One complication is that we may need to turn off the raster magnets at the straight-ahead dump to use the last viewer. The overall procedure is not yet finalized, but it is in process of being developed.
- Multimonitor works, but the uncertainty in the measurements is presently too large (±50%) for the data to be useful. More refinement is needed.
- M_55 cavities work as bunch-length monitors; the associated software has been checked.
- Zero phasing works, but we need good steering through the cryomodule cavities to get good data. There is presently no approach that is reproducible operator to operator at the needed accuracy, so work on the technique continues.
- Happek (University of Georgia) is scheduled to be here from
10-13 May 98. We will install his hardware during day shift Monday
and test it during swing and owl shifts Monday and Tuesday nights
As a general remark, phase-space parameters that have been measured
thus far generally fall within a factor 2-3 (far less than an
order of magnitude!) from those required for lasing. Of course,
if all parameters were double the requirement, no lasing could
happen. The point is that the data continue to suggest we're
within reach but need to continue fine-tuning the beam before
installing the wiggler.
Significant progress was made toward setting up the FEL optical
systems. Optical transport diagnostics were completed, aligned
and put under vacuum, and they were used to take a surrogate HeNe
beam from the FEL to the diagnostics table in the optical control
room. Patch panels to the optical control room were installed
and fully wired. Fiducialization of the first-light mirrors was
completed. The mirror-control software for save/restore was verified.
The cavity mirror alignment as viewed through position sensors
drifted by less than 10 µrad over several days, well within
Measurements of drive-laser noise were recently made. The principal observations are:
The drive laser performs best when running on its internal rf system. Performance is comparable to the required value for (FEL) lasing: <3% p-p at 10 Hz > f > 50 kHz, but worse than measured when the laser was new (1.3% p-p on 3/30/95).
When running on external rf, amplitude noise is higher. Closing
the phase-locked loop results in amplitude noise of order 18-19%.
We are beginning to prepare a plan to correct this problem.
Over the next two weeks the basic program is to do orbit and phase-space optimization during day shifts, cw runs during swing shifts, and beam diagnostics/beam quality measurements during owl shifts. Progress from each shift will feed into the other shifts as together we iterate toward a usable lasing setup.